# Rethinking Robustness in Machine Learning: A Posterior Agreement Approach

### v1

@@ 4.2 Domain Generalization
While our metric has been primarily intended for evaluation, in this section, we investigate whether PA can also effectively guide model selection. To that end, we now examine a setting where the inductive bias is deliberately “poisoned” by shortcut opportunities (SO), i.e., spurious correlations between the predicted factor $F_{P}$ and the learning factor $F_{L}$. As illustrated in Figure 12 (Appendix E), we systematically control these co-occurrences to create zero or partial Generalization Opportunities (GO). Under Zero-GO (ZGO), each instance of $F_{P}$ is exclusively paired with one instance of $F_{L}$, strongly encouraging models to overfit to the shortcut. Partial GOs ($1/2/3$-CGO) break some of these exclusive pairings, while Zero-SO (ZSO) allows all factor combinations. For more details on the experiment setup, refer to Appendix E.
@@
Acc. $\Delta\textbf{Acc.}$. Acc. $\Delta\textbf{Acc.}$. Acc. $\Delta\textbf{Acc.}$. Acc. $\Delta\textbf{Acc.}$. Acc. $\Delta\textbf{Acc.}$.
@@
Table 1: Test performance under increasing levels of shift for models selected through different configurations of factor co-occurrence in the hue-based learning factor experiment. Specifically, the performance of models selected through validation accuracy (Acc) and the difference between accuracy-based and PA-based selection ($\Delta$Acc) are reported. PA is able to select models that perform better than the accuracy-selected model in most cases.
@@ 5 Discussion
The PA metric is an attempt at assessing robustness under an epistemologically grounded approach. Such an assessment is agnostic of the underlying data generation process and even the nature of the data itself, therefore it is versatile and can be applied to explain diverse scenarios with a single theory, as illustrated in the experimental study. Additionally, PA is general in its scope, as it only requires the (probabilistic) outputs of the model. Our proposed version requires only the logits of a model [^6] N.B. The model itself does not need to be differentiable, as we do not optimize its parameters. to optimize a single parameter $\beta$, however, variants are possible (cf. Buhmann (2010); Buhmann et al. (2018)). By relying on the MEP, PA provides an estimate of model’s robustness which is as neutral as possible with respect to missing information, while sticking to the available observations. In addition to that, its Bayesian perspective, allows us to give a robustness estimation of the entire hypothesis set and not just the specific model.
The experimental results show that PA evaluates models at a finer level, providing higher discriminability and consistency in covariate shift settings than accuracy-based metrics. The discriminative power of $\text{AFR}_{T}$ relies, instead, solely on test performance and does not fulfill the desired properties of a robustness metric. Additionally, by relying on predictive confidence rather than preditions, PA favors classifiers that exhibit desirable inductive biases, and grasps sensitivity to the nature of the shift at an even greater detail than in standard analyses.
Our proposed version of PA has, however, limitations. The theoretical framework has been developed by considering only the case of a finite hypothesis set. While this does not pose a severe drawback, and the experimental results attest to the validity of this approach, a characterization of PA for continuous hypothesis sets would be much more general in its scope. Additionally, we have assumed a uniform probability $p(c)$, to obtain a more tractable formula for the metric, ignoring the complexity of the hypothesis set. Adding this information would improve the robustness estimation process, allowing for more fine-grained discrimination between models. In future, we plan to extend some of the presented theoretical results to more general kernels.
@@ 6 Conclusions
In this work, we have set some desiderata for designing a metric measuring robustness against forms of covariate shift such as adversarial noise and domain generalization. Following theoretically grounded thinking, we proposed PA, a metric derived from the Posterior Agreement framework. We have conducted a comparison against commonly used accuracy metrics, presenting analyses based on novel aspects of the shift. Our study shows that PA is consistent with the desired robustness properties. Additionally, by estimating the overlap between posteriors over the hypothesis space, PA provides a more fine-grained evaluation and a better discriminative power than hard-counting metrics, such as accuracy. In conclusion, our work lays the basis for a more sound model robustness assessment, in the PA sense.
@@ A.3 Proof of Theorem 3
$p_{i}(j\mid X)=\frac{\exp(\beta R(j,x_{i}))}{\sum_{k=1}^{K}\exp(\beta R(k,x_{k}))}.$ (27)
@@
Let us define $R(j,x^{\prime})=R_{j}^{\prime}$ and $\exp(\beta R(j,x^{\prime}))=\exp(\beta R_{j}^{\prime})=e_{j}^{\prime}$, with $x^{\prime}\in X^{\prime}$. Similar notation is used for $x^{\prime\!\>\!\prime}\in X^{\prime\!\>\!\prime}$. $f(\beta)$ can be therefore rewritten as
@@
$\frac{d}{d\beta}\lambda_{1}(\beta)=\frac{\sum_{j=1}^{K}(R_{j}^{\prime}+R_{j}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}}{\sum_{k=1}^{K}e_{k}^{\prime}e_{k}^{\prime\!\>\!\prime}}.$ (30)
@@
$\frac{d^{2}}{d\beta^{2}}\lambda_{1}(\beta)=\frac{\sum_{j=1}^{K}(R_{j}^{\prime}+R_{j}^{\prime\!\>\!\prime})^{2}e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}}{\sum_{k=1}^{K}e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}}-\left(\frac{\sum_{j=1}^{K}(R_{j}^{\prime}+R_{j}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}}{\sum_{k=1}^{K}e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}}\right)^{2}.$ (31)
@@
$\frac{d^{2}}{d\beta^{2}}\lambda_{1}(\beta)>0\iff\left(\sum_{k=1}^{K}e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}\right)\left(\sum_{j=1}^{K}(R_{j}^{\prime}+R_{j}^{\prime\!\>\!\prime})^{2}e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}\right)-\left(\sum_{j=1}^{K}(R_{j}^{\prime}+R_{j}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}\right)^{2}>0$ (32)
@@
$\sum_{k=1}^{K}\sum_{j=1}^{K}(R_{j}^{\prime}+R_{j}^{\prime\!\>\!\prime})^{2}e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}e_{k}^{\prime}e_{k}^{\prime\!\>\!\prime}-\sum_{k=1}^{K}\sum_{j=1}^{K}(R_{j}^{\prime}+R_{j}^{\prime\!\>\!\prime})(R_{k}^{\prime}+R_{k}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}e_{k}^{\prime}e_{k}^{\prime\!\>\!\prime}>0$ (33)
$\iff\sum_{k=1}^{K}\sum_{j=1}^{K}[(R_{j}^{\prime}+R_{j}^{\prime\!\>\!\prime})-(R_{k}^{\prime}+R_{k}^{\prime\!\>\!\prime})](R_{j}^{\prime}+R_{j}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}e_{k}^{\prime}e_{k}^{\prime\!\>\!\prime}>0$ (34)
Let $\Delta_{(jj),(kk)}=(R_{j}^{\prime}+R_{j}^{\prime\!\>\!\prime})-(R_{k}^{\prime}+R_{k}^{\prime\!\>\!\prime})$ define the difference in the cost attributed to reference class $j$ and the cost attributed to class $k$, accumulated over $X^{\prime},X^{\prime\!\>\!\prime}$, and $E_{jk}=e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}e_{k}^{\prime}e_{k}^{\prime\!\>\!\prime}$.
@@
$\sum_{k=1}^{K}\sum_{j=1}^{K}[(R_{j}^{\prime}+R_{j}^{\prime\!\>\!\prime})-(R^{\prime}_{k}+R_{k}^{\prime\!\>\!\prime})](R_{j}^{\prime}+R_{j}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}e_{k}^{\prime}e_{k}^{\prime\!\>\!\prime}=\sum_{k=1}^{K}\sum_{j=1}^{K}(R_{j}^{\prime}+R_{j}^{\prime\!\>\!\prime})E_{jk}\Delta_{(jj),(kk)}=\sum_{k=1}^{K}\sum_{j=1}^{K}S_{(jj),(kk)}$ (35)
Note that $\Delta_{(jj),(jj)}=0\implies S_{(jj),(jj)}=0$. Moreover, $\Delta_{(jj),(kk)}=-\Delta_{(kk),(jj)}$ and $E_{jk}=E_{kj}$. Then, the previous term can be rewritten as
$\sum_{j=1}^{K}\sum_{k<j}^{K}S_{(jj),(kk)}+S_{(kk),(jj)}=\sum_{j=1}^{K}\sum_{k<j}^{K}(R_{j}^{\prime}+R_{j}^{\prime\!\>\!\prime})E_{jk}\Delta_{(jj),(kk)}+(R_{k}^{\prime}+R_{k}^{\prime\!\>\!\prime})E_{kj}\Delta_{(kk),(jj)}$ (36)
$=\sum_{j=1}^{K}\sum_{k<j}^{K}E_{jk}\Delta_{(jj),(kk)}[(R_{j}^{\prime}+R_{j}^{\prime\!\>\!\prime})-(R_{k}^{\prime}+R_{k}^{\prime\!\>\!\prime})]=\sum_{j=1}^{K}\sum_{k<j}^{K}E_{jk}\Delta_{(jj),(kk)}^{2}$ (37)
@@
$\frac{d}{d\beta}\lambda_{2}(\beta)=\frac{\sum_{k=1}^{K}\sum_{j=1}^{K}(R_{j}^{\prime}+R_{k}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}}{\sum_{k=1}^{K}\sum_{j=1}^{K}e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}}$ (39)
$\frac{d^{2}}{d\beta^{2}}\lambda_{2}(\beta)=\frac{\sum_{k=1}^{K}\sum_{j=1}^{K}(R_{j}^{\prime}+R_{k}^{\prime\!\>\!\prime})^{2}e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}}{\sum_{k=1}^{K}\sum_{j=1}^{K}e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}}-\left(\frac{\sum_{k=1}^{K}\sum_{j=1}^{K}(R_{j}^{\prime}+R_{k}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}}{\sum_{k=1}^{K}\sum_{j=1}^{K}e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}}\right)^{2}>0$ (40)
$\iff\left(\sum_{k=1}^{K}\sum_{j=1}^{K}e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}\right)\left(\sum_{k=1}^{K}\sum_{j=1}^{K}(R_{j}^{\prime}+R_{k}^{\prime\!\>\!\prime})^{2}e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}\right)-\left(\sum_{k=1}^{K}\sum_{j=1}^{K}(R_{j}^{\prime}+R_{k}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}\right)^{2}>0$ (41)
$\iff\sum_{k=1}^{K}\sum_{q=1}^{K}\sum_{j=1}^{K}\sum_{i=1}^{K}(R_{j}^{\prime}+R_{k}^{\prime\!\>\!\prime})^{2}e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}e_{i}^{\prime}-(R_{j}^{\prime}+R_{k}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}(R_{i}^{\prime}+R_{q}^{\prime\!\>\!\prime})e_{i}^{\prime}e_{i}^{\prime\!\>\!\prime}>0$ (42)
$\iff\sum_{k=1}^{K}\sum_{q=1}^{K}\sum_{j=1}^{K}\sum_{i=1}^{K}(R_{j}^{\prime}+R_{k}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}e_{i}^{\prime}e_{i}^{\prime\!\>\!\prime}[(R_{j}^{\prime}+R_{k}^{\prime\!\>\!\prime})-(R_{i}^{\prime}+R_{q}^{\prime\!\>\!\prime})]>0$ (43)
Similarly, as we did before, we define $e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}e_{i}^{\prime}e_{i}^{\prime\!\>\!\prime}=E_{(jk),(iq)}=E_{(ik),(jq)}=E_{(jq),(ik)}=E_{(iq),(jk)}$ and $\Delta_{(jk),(iq)}=(R_{j}^{\prime}-R_{i}^{\prime})+(R_{k}^{\prime\!\>\!\prime}-R_{q}^{\prime\!\>\!\prime})=-\Delta_{(iq),(jk)}$. Then,
$\frac{d^{2}}{d^{2}\beta}\lambda_{2}(\beta)=\sum_{k=1}^{K}\sum_{q=1}^{K}\sum_{j=1}^{K}\sum_{i=1}^{K}S_{(jk),(iq)}=\sum_{k=1}^{K}\sum_{q=1}^{K}\sum_{j=1}^{K}\sum_{i=1}^{K}(R_{j}^{\prime}+R_{k}^{\prime\!\>\!\prime})E_{(jk),(iq)}\Delta_{(jk),(iq)}$ (44)
@@
$\sum_{k=1}^{K}\sum_{q<k}^{K}\sum_{j=1}^{K}\sum_{i<j}^{K}S_{(jk),(iq)}+S_{(iq),(jk)}$ (45)
$=\sum_{k=1}^{K}\sum_{q<k}^{K}\sum_{j=1}^{K}\sum_{i<j}^{K}(R_{j}^{\prime}+R_{k}^{\prime\!\>\!\prime})E_{(jk),(iq)}\Delta_{(jk),(iq)}+(R_{i}^{\prime}+R_{q}^{\prime\!\>\!\prime})E_{(iq),(jk)}\Delta_{(iq),(jk)}$ (46)
$=\sum_{k=1}^{K}\sum_{q<k}^{K}\sum_{j=1}^{K}\sum_{i<j}^{K}E_{(jk),(iq)}\Delta_{(jk),(iq)}[(R_{j}^{\prime}+R_{k}^{\prime\!\>\!\prime})-(R_{i}^{\prime}+R_{q}^{\prime\!\>\!\prime})]$ (47)

### v2

@@ 4.2 Domain Generalization
While our measure has been primarily intended for evaluation, in this section, we investigate whether PA can also effectively guide model selection. To that end, we now examine a setting where the inductive bias is deliberately “poisoned” by shortcut opportunities (SO), i.e., spurious correlations between the predicted factor $F_{P}$ and the learning factor $F_{L}$. We train a ResNet5050 architecture from scratch using datasets $X^{\prime}$ and $X^{\prime\prime}$, which differ only in the hue factor. These data sets are resampled for each shortcut-opportunity configuration, where spurious correlations between the predicted factor $F_{P}$ and the learning factor $F_{L}$ are systematically controlled. As illustrated in Figure 12 (and in more detail Appendix E), we systematically control these co-occurrences to create zero or partial Generalization Opportunities (GO). Under Zero-GO (ZGO), each instance of $F_{P}$ is exclusively paired with one instance of $F_{L}$, strongly encouraging models to overfit to the shortcut. Partial GOs ($1/2/3$-CGO) break some of these exclusive pairings, while Zero-SO (ZSO) allows all factor combinations. Generalization performance is then evaluated on progressively shifted test datasets, equivalent to those used in our previous experiment. Epoch-wise model selection is performed by either maximizing accuracy or maximizing PA on the validation sets, and the difference in generalization performance between these criteria is reported. For more details on the experiment setup, including the precise train/validation configurations, refer to Appendix E.
@@
Acc. $\Delta\textbf{Acc.}$ Acc. $\Delta\textbf{Acc.}$ Acc. $\Delta\textbf{Acc.}$ Acc. $\Delta\textbf{Acc.}$ Acc. $\Delta\textbf{Acc.}$
@@
[…]
@@ 5 Discussion
The PA measure is an attempt at assessing robustness under an epistemologically grounded approach. Such an assessment is agnostic of the underlying data generation process and even the nature of the data itself, therefore, it is versatile and can be applied to explain diverse scenarios with a single theory, as illustrated in the experimental study. Additionally, PA is general in its scope. Our proposed version requires only the probabilistic outputs of the model, e.g., the logits of a model, and no supervision. The model itself does not need to be differentiable, as we do not optimize its parameters.
By relying on the MEP, PA provides an estimate of a model’s robustness which is as neutral as possible with respect to missing information. The measure relies on predictive confidence rather than predictions, thus providing a more fine-grained evaluation and a better discriminative assessment of model robustness. Indeed, the overlap between posteriors over the hypothesis space is indicative of the nature of the shift at an even greater detail than in standard analyses. Instead, test performance measures, such as $\text{AFR}_{T}$ rely solely on hard-counting of “successes”, therefore, they do not fulfill the desired properties of a robustness measure and are more prone to inconsistencies in the rankings.
In particular, in the adversarial learning setting, PA provides higher discriminability and consistency across the various covariate shifts, which is only partially matched by AFR in highly perturbed settings. Additionally, the analysis of the $\beta$ parameter provides further insights into the robustness of the models. In the domain generalization setting, PA further highlights that the robustness of the tested models decreases only with the first shifts, then it stabilizes. $\text{AFR}_{T}$, instead, keeps decreasing, showing once again that robustness and performance are two distinct concepts. In this setting, PA also proves to be very efficient in performing model selection.
Our proposed version of PA has, however, limitations. The theoretical framework has been developed by considering only the case of a finite hypothesis set. While this does not pose a severe drawback, and the experimental results attest to the validity of this approach, a characterization of PA for continuous hypothesis sets would be much more general in its scope. Additionally, we have assumed a uniform probability $p(c)$, to obtain a more tractable formula for the measure, ignoring the complexity of the hypothesis set. Adding this information would improve the robustness estimation process, allowing for more fine-grained discrimination between models. In the future, we plan to extend some of the presented theoretical results to more general kernels.
@@ 6 Conclusions
In this paper, we have set some desiderata for designing a measure measuring robustness against forms of covariate shift such as adversarial noise and domain generalization. Following theoretically grounded thinking, we proposed PA, a measure derived from the Posterior Agreement framework. We have conducted a comparison against commonly used accuracy measures, presenting analyses based on novel aspects of the shift. Our study shows that PA (i) provides a consistent and reliable robustness scoring, (ii) provides further details to the analysis that would be otherwise overlooked by accuracy measures and (iii) is also an efficient measure for model selection, while requiring no supervision. In conclusion, our work lays the basis for a more sound model robustness assessment, in the PA sense.
@@ A.3 Proof of Theorem 3
$p_{i}(j\mid X)=\frac{\exp(\beta\mathcal{S}(j,x_{i}))}{\sum_{k=1}^{K}\exp(\beta\mathcal{S}(k,x_{k}))}.$ (29)
@@
Let us define $\mathcal{S}(j,x^{\prime})=\mathcal{S}_{j}^{\prime}$ and $\exp(\beta\mathcal{S}(j,x^{\prime}))=\exp(\beta\mathcal{S}_{j}^{\prime})=e_{j}^{\prime}$, with $x^{\prime}\in X^{\prime}$. Similar notation is used for $x^{\prime\!\>\!\prime}\in X^{\prime\!\>\!\prime}$. $f(\beta)$ can be therefore rewritten as
@@
$\frac{d}{d\beta}\lambda_{1}(\beta)=\frac{\sum_{j=1}^{K}(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{j}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}}{\sum_{k=1}^{K}e_{k}^{\prime}e_{k}^{\prime\!\>\!\prime}}.$ (32)
@@
$\frac{d^{2}}{d\beta^{2}}\lambda_{1}(\beta)=\frac{\sum_{j=1}^{K}(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{j}^{\prime\!\>\!\prime})^{2}e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}}{\sum_{k=1}^{K}e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}}-\left(\frac{\sum_{j=1}^{K}(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{j}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}}{\sum_{k=1}^{K}e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}}\right)^{2}.$ (33)
@@
$\frac{d^{2}}{d\beta^{2}}\lambda_{1}(\beta)>0\iff\left(\sum_{k=1}^{K}e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}\right)\left(\sum_{j=1}^{K}(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{j}^{\prime\!\>\!\prime})^{2}e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}\right)-\left(\sum_{j=1}^{K}(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{j}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}\right)^{2}>0$ (34)
@@
$\sum_{k=1}^{K}\sum_{j=1}^{K}(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{j}^{\prime\!\>\!\prime})^{2}e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}e_{k}^{\prime}e_{k}^{\prime\!\>\!\prime}-\sum_{k=1}^{K}\sum_{j=1}^{K}(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{j}^{\prime\!\>\!\prime})(\mathcal{S}_{k}^{\prime}+\mathcal{S}_{k}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}e_{k}^{\prime}e_{k}^{\prime\!\>\!\prime}>0$ (35)
$\iff\sum_{k=1}^{K}\sum_{j=1}^{K}[(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{j}^{\prime\!\>\!\prime})-(\mathcal{S}_{k}^{\prime}+\mathcal{S}_{k}^{\prime\!\>\!\prime})](\mathcal{S}_{j}^{\prime}+\mathcal{S}_{j}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}e_{k}^{\prime}e_{k}^{\prime\!\>\!\prime}>0$ (36)
Let $\Delta_{(jj),(kk)}=(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{j}^{\prime\!\>\!\prime})-(\mathcal{S}_{k}^{\prime}+\mathcal{S}_{k}^{\prime\!\>\!\prime})$ define the difference in the cost attributed to reference class $j$ and the cost attributed to class $k$, accumulated over $X^{\prime},X^{\prime\!\>\!\prime}$, and $E_{jk}=e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}e_{k}^{\prime}e_{k}^{\prime\!\>\!\prime}$.
@@
$\sum_{k=1}^{K}\sum_{j=1}^{K}[(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{j}^{\prime\!\>\!\prime})-(\mathcal{S}_{k}^{\prime}+\mathcal{S}_{k}^{\prime\!\>\!\prime})](\mathcal{S}_{j}^{\prime}+\mathcal{S}_{j}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{j}^{\prime\!\>\!\prime}e_{k}^{\prime}e_{k}^{\prime\!\>\!\prime}=\sum_{k=1}^{K}\sum_{j=1}^{K}(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{j}^{\prime\!\>\!\prime})E_{jk}\Delta_{(jj),(kk)}=\sum_{k=1}^{K}\sum_{j=1}^{K}G_{(jj),(kk)}$ (37)
Note that $\Delta_{(jj),(jj)}=0\implies G_{(jj),(jj)}=0$. Moreover, $\Delta_{(jj),(kk)}=-\Delta_{(kk),(jj)}$ and $E_{jk}=E_{kj}$. Then, the previous term can be rewritten as
$\sum_{j=1}^{K}\sum_{k<j}^{K}G_{(jj),(kk)}+G_{(kk),(jj)}=\sum_{j=1}^{K}\sum_{k<j}^{K}(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{j}^{\prime\!\>\!\prime})E_{jk}\Delta_{(jj),(kk)}+(\mathcal{S}_{k}^{\prime}+\mathcal{S}_{k}^{\prime\!\>\!\prime})E_{kj}\Delta_{(kk),(jj)}$ (38)
$=\sum_{j=1}^{K}\sum_{k<j}^{K}E_{jk}\Delta_{(jj),(kk)}[(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{j}^{\prime\!\>\!\prime})-(\mathcal{S}_{k}^{\prime}+\mathcal{S}_{k}^{\prime\!\>\!\prime})]=\sum_{j=1}^{K}\sum_{k<j}^{K}E_{jk}\Delta_{(jj),(kk)}^{2}$ (39)
@@
$\frac{d}{d\beta}\lambda_{2}(\beta)=\frac{\sum_{k=1}^{K}\sum_{j=1}^{K}(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{k}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}}{\sum_{k=1}^{K}\sum_{j=1}^{K}e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}}$ (41)
$\frac{d^{2}}{d\beta^{2}}\lambda_{2}(\beta)=\frac{\sum_{k=1}^{K}\sum_{j=1}^{K}(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{k}^{\prime\!\>\!\prime})^{2}e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}}{\sum_{k=1}^{K}\sum_{j=1}^{K}e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}}-\left(\frac{\sum_{k=1}^{K}\sum_{j=1}^{K}(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{k}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}}{\sum_{k=1}^{K}\sum_{j=1}^{K}e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}}\right)^{2}>0$ (42)
$\iff\left(\sum_{k=1}^{K}\sum_{j=1}^{K}e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}\right)\left(\sum_{k=1}^{K}\sum_{j=1}^{K}(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{k}^{\prime\!\>\!\prime})^{2}e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}\right)-\left(\sum_{k=1}^{K}\sum_{j=1}^{K}(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{k}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}\right)^{2}>0$ (43)
$\iff\sum_{k=1}^{K}\sum_{q=1}^{K}\sum_{j=1}^{K}\sum_{i=1}^{K}(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{k}^{\prime\!\>\!\prime})^{2}e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}e_{i}^{\prime}-(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{k}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}(\mathcal{S}_{i}^{\prime}+\mathcal{S}_{q}^{\prime\!\>\!\prime})e_{i}^{\prime}e_{i}^{\prime\!\>\!\prime}>0$ (44)
$\iff\sum_{k=1}^{K}\sum_{q=1}^{K}\sum_{j=1}^{K}\sum_{i=1}^{K}(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{k}^{\prime\!\>\!\prime})e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}e_{i}^{\prime}e_{i}^{\prime\!\>\!\prime}[(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{k}^{\prime\!\>\!\prime})-(\mathcal{S}_{i}^{\prime}+\mathcal{S}_{q}^{\prime\!\>\!\prime})]>0$ (45)
Similarly, as we did before, we define $e_{j}^{\prime}e_{k}^{\prime\!\>\!\prime}e_{i}^{\prime}e_{i}^{\prime\!\>\!\prime}=E_{(jk),(iq)}=E_{(ik),(jq)}=E_{(jq),(ik)}=E_{(iq),(jk)}$ and $\Delta_{(jk),(iq)}=(\mathcal{S}_{j}^{\prime}-\mathcal{S}_{i}^{\prime})+(\mathcal{S}_{k}^{\prime\!\>\!\prime}-\mathcal{S}_{q}^{\prime\!\>\!\prime})=-\Delta_{(iq),(jk)}$. Then,
$\frac{d^{2}}{d^{2}\beta}\lambda_{2}(\beta)=\sum_{k=1}^{K}\sum_{q=1}^{K}\sum_{j=1}^{K}\sum_{i=1}^{K}G_{(jk),(iq)}=\sum_{k=1}^{K}\sum_{q=1}^{K}\sum_{j=1}^{K}\sum_{i=1}^{K}(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{k}^{\prime\!\>\!\prime})E_{(jk),(iq)}\Delta_{(jk),(iq)}$ (46)
@@
$\sum_{k=1}^{K}\sum_{q<k}^{K}\sum_{j=1}^{K}\sum_{i<j}^{K}G_{(jk),(iq)}+G_{(iq),(jk)}$ (47)
$=\sum_{k=1}^{K}\sum_{q<k}^{K}\sum_{j=1}^{K}\sum_{i<j}^{K}(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{k}^{\prime\!\>\!\prime})E_{(jk),(iq)}\Delta_{(jk),(iq)}+(\mathcal{S}_{i}^{\prime}+\mathcal{S}_{q}^{\prime\!\>\!\prime})E_{(iq),(jk)}\Delta_{(iq),(jk)}$ (48)
$=\sum_{k=1}^{K}\sum_{q<k}^{K}\sum_{j=1}^{K}\sum_{i<j}^{K}E_{(jk),(iq)}\Delta_{(jk),(iq)}[(\mathcal{S}_{j}^{\prime}+\mathcal{S}_{k}^{\prime\!\>\!\prime})-(\mathcal{S}_{i}^{\prime}+\mathcal{S}_{q}^{\prime\!\>\!\prime})]$ (49)